# Fast, accurate and lightweight sequential simulation-based inference using Gaussian locally linear mappings

### 12 Mar 2024 v1

@@ 4 SeMPLE: sequential learning of likelihoods and posteriors
We have shown how GLLiM can provide amortized estimations of both the likelihood and the posterior distribution at no additional cost, given training data. This GLLiM feature is exploited in our proposed SeMPLE procedure, to obtain a sequence of increasingly improved samplers, to more accurately and more efficiently sample from the targeted posterior of observed data $\bm{y}_{o}$. Similarly to SNL (Papamakarios et al., 2019), given a likelihood approximation $q_{\tilde{\bm{\phi}}}(\bm{y}_{o}\,|\,\bm{\theta})$, SeMPLE targets an approximate posterior $\tilde{p}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ with posterior sampling being carried out using a Markov Chain Monte Carlo (MCMC) procedure. Typically, the MCMC scheme used in SNL is slice sampling (Neal, 2003), which has the advantage to be tuning-free but is likely to have difficulties in exploring high dimensional posteriors with complex multimodalities. As a matter of fact, Papamakarios et al. (2019) recommend SNL “at most for tens of parameters that do not have pathologically strong correlations”. In SeMPLE, since GLLiM is able to learn a closed-form approximation of the posterior distribution (in addition to a likelihood function), such posterior is used as a proposal distribution within a Metropolis-Hastings (MH) algorithm. SeMPLE, by using an efficient EM procedure to fit a flexible proposal distribution, is expected to benefit from a MH step that better guides simulations towards more likely parameter values. To ease the comparison with SNL, the latter is recalled in Supplementary Section 4, while SeMPLE is summarized in Algorithms 1-2.
@@
Algorithm 1 shows that the targeted approximate posterior for $\bm{\theta}$ is a function denoted with $\hat{p}_{r}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ which is sequentially obtained, where $r=0,...,R$ denotes the SeMPLE rounds for the current approximation [^2] We use the term round instead of the most common iteration, as the latter could be confused with the MCMC iterations that are run within a given SeMPLE round.. In the following we denote by $\bm{\theta}^{(r)}$ and $\bm{y}^{(r)}$ a parameter and an observation simulated in round $r$, respectively. At first ($r=0$) $\hat{p}_{0}(\bm{\theta})$ is initialised at the prior $p(\bm{\theta})$. An initial batch of $N$ pairs $\mathcal{D}_{0}=\{\bm{\theta}_{n}^{(0)},\bm{y}_{n}^{(0)}\}_{n=1}^{N}$ produced from the prior-predictive distribution (steps 4-5) constitutes the starting training data for GLLiM, and the parameters $\tilde{\phi}_{0}$ and ${\phi}_{0}$ are obtained via EM within GLLiM to produce the initial surrogate likelihood and posterior $q_{\tilde{\bm{\phi}}_{0}}(\bm{y}\,|\,\bm{\theta})$ and $q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y})$ respectively (steps 9-10). The initial surrogate posterior becomes the current target (step 11) $\tilde{p}_{1}(\bm{\theta})\equiv q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y}_{o})$. In the next iteration ($r=1$), the latter can be sampled from without MH (step 4) since $\tilde{p}_{1}(\bm{\theta})$ is defined as a Gaussian mixture, from which it is trivial to sample $N$ parameters $\{\bm{\theta}_{n}^{(1)}\}_{n=1}^{N}$. These $N$ draws are then plugged into the model simulator to produce $\bm{y}_{n}^{(1)}\sim p(\bm{y}\,|\,\bm{\theta}_{n}^{(1)})$ and define $\mathcal{D}_{1}=\{\bm{\theta}_{n}^{(1)},y_{n}^{(1)}\}_{n=1}^{N}$ as the new training data (we discard $\mathcal{D}_{0}$ for future iterations as it is deemed too uninformative for training). GLLiM is then fitted to $\mathcal{D}_{1}$ and the corresponding surrogate likelihood and surrogate posterior are obtained. From this step on, the surrogate posterior is employed as a proposal distribution within a tuning-free MH algorithm (Algorithm 2) since, for an arbitrary prior $p(\bm{\theta})$, when $r>1$ we do not explicitly know which distribution is proportional to the product $p(\bm{\theta})q_{\tilde{\phi}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})$ (for $r=1$ we know by construction that the targeted posterior is a Gaussian mixture), except for the special case of Gaussian and Uniform priors (see Papamakarios and Murray, 2016). In Algorithm 2, the acceptance probability of a proposed $\bm{\theta}^{*}$ is (for $r>1$)
@@
Importantly, our MCMC procedure is tuning-free. We use an independence sampler (Robert and Casella, 2004), relying on GLLiM to provide an informative proposal function. This proposal function is trained on draws accepted in previous rounds which are, reasonably, more spread than the currently targeted posterior. In addition, being a mixture model, it is suitable to deal with multimodalities in the target. This greatly simplifies the construction of SeMPLE as no complex tuning is necessary for the MCMC sampler. While $K$ could in principle be set arbitrarily large, in practice this may slow the EM algorithm convergence. In SeMPLE, the value of $K$ is allowed to decrease between iterations in the update $K$ step lines 9 and 14 in Algorithm 1. Mixture components having probabilities $\tilde{\eta}_{k}$ smaller than a threshold are discarded. In most cases we find that setting a starting value of $K=20$ or 30 produces good results, with a threshold for $\tilde{\eta}_{k}$ typically set to 0.005 (multiple hyperboloid, Ornstein-Uhlenbeck, Lotka-Volterra models), 0.03 (Bernoulli GLM model) or simply 0 for no deletion (Two Moons model). This is to prevent having too many irrelevant components, that could be problematic when fitting the mixture to training data produced via MCMC, as the latter may not manage to visit tiny modes. It is also possible to choose the starting value of $K$ in a principled way. In Deleforge et al. (2014), the Bayesian Information Criterion (BIC) is used to select $K$, as we do in our examples, see Supplementary Section 2. An appropriate value for $K$ can be learned by repeatedly computing the BIC on realizations of the prior-predictive distribution, before SeMPLE is started. This is an amortized procedure, as the selected values can be reused regardless the specific observed data $\bm{y}_{o}$. The impact on the overall computational budget is negligible. We have also experimented with a version of SeMPLE that learns a corrected posterior (see the optional part in line 16 of Algorithm 1) using the correction formula from Papamakarios and Murray (2016). However, we found that learning a surrogate likelihood was much more beneficial, providing better inference coupled with a higher acceptance rate. Notice that none of our reported results use the optional corrected-posterior approach.
@@ 5 Numerical illustrations
We illustrate SeMPLE on several simulation studies, see also the Supplementary Material for extra results and additional models. SeMPLE is written in R and uses the GLLiM implementation from the xLLiM package (Perthame et al., 2022). Comparisons with neural-based likelihood and posterior estimation methods (SNL and SNPE-C) use the Python implementations in the SBIBM package for SBI benchmarking (Lueckmann et al., 2021), and architecture specifications for SBIBM are in Supplementary Section 1. To compare approximate posteriors with a suitable reference posterior, we consider the classifier two-samples test (C2ST), as implemented in SBIBM, and Wasserstein distances via the POT Python package (Flamary et al., 2021). C2ST varies in the interval [0.5,1], the lower the better. It equals 0.5 when samples from two distributions appear statistically indistinguishable and equals 1 when the classifier is able to perfectly separate samples as belonging to two different distributions. We also report runtimes in figures and resource requirements in Table 1. For SeMPLE, the time required to select $K$ via BIC is reported separately and not taken into account in the figures below. For some case studies, such as Lotka-Volterra, we merely picked a starting value of $K$ without running the BIC procedure. Adding this overhead time does not change the overall conclusion that SeMPLE is much more time-efficient than the other methods. Moreover, the selection of $K$ is an amortized procedure, hence it does not need to be repeated for different observations of the same model.
@@
In the next two sections 5.1 and 5.2, we illustrate results from two models having multimodal posteriors, in section 5.3 we summarise the Lotka-Volterra experiment, and in section 5.5 and Table 2 we summarise results from all other experiments detailed in Supplementary Material. For comparison, all methods are given the same budget of model simulations. For SNL and SNPE-C the SBIBM default of distributing these simulations uniformly across $R=10$ rounds is used, but to ease comparison we also run SNL and SNPE-C with the same number of rounds used for SeMPLE, which is $R=4$, except for the Bernoulli GLM model where $R=2$ suffices. In fact, we observe that SeMPLE does not need as many sequential steps as the other methods. It may be argued that SNL could perform better if a larger number of MCMC iterations was used, at the cost of a smaller number of rounds $R$. Results in Supplementary Material show that both SNL and SNPE-C perform worse under this alternative configuration. Regarding SeMPLE, when the computational budget limits the number of model simulations, we found it beneficial to keep the number of rounds $R$ between 2 and 3, where $R=2$ implies that we have exactly one round where MCMC is used, and this is typically already enough for satisfying inference, though usually an additional round with MCMC (hence $R=3$) produces further refined inference. In our work we show results up to $R=4$, but in general, four rounds are not recommended. The algorithm performance relies first on the GLLiM ability to learn reasonably accurate likelihood and posterior approximations and then on the MCMC sampler to explore the parameter space. Both steps require a large enough number of model simulations, which depends mainly on the model dimension.
@@ 5.1 Two Moons
The Two Moons model is a standard benchmark example in SBI (Greenberg et al., 2019; Lueckmann et al., 2021; Picchini and Tamborrino, 2022) to illustrate how algorithms deal with multimodality and the local crescent shapes of the posterior distribution. The model itself is defined in Supplementary Section 5. Our experiments are run on the same 10 datasets provided in SBIBM, each of them being a vector $\bm{y}_{0}$ of length two. We use $K=30$, a value we determine as described in Supplementary Material (this determination only requires 90 seconds). We run SeMPLE, SNL and SNPE-C algorithms for a total budget of $10^{4}$ model simulations, distributed across inference rouces, and we tested both $R=4$ and $R=10$ rounds. The obtained samples are then compared with the corresponding posterior reference samples (obtained via MCMC, since the likelihood is tractable) provided in SBIBM. The performance metric results are averaged over the 10 different data sets. Figure 2 shows C2ST and runtimes values. SeMPLE provides the best approximation while SNL performs the worst. The comparison with SNL is particularly interesting as while both SNL and SeMPLE incorporate a tuning-free MCMC step, SeMPLE is able to visit both modes with an acceptance rate of 60-70%, while SNL often struggles in efficiently exploring a multimodal posterior as demonstrated by the higher C2ST values. Moreover, in this simple example the runtime for SeMPLE is about half the SNPE-C runtime, and is about three times lower than the SNL runtime. In other examples (e.g. Figure 3) the speedup provided by SeMPLE is even more dramatic. All these benefits from SeMPLE come with a very small impact on the memory usage, see Section 5.4 for details.
@@ 5.2 Multiple hyperboloid model
The multiple hyperboloid model (Forbes et al., 2022) is constructed from a sound source localization problem in audio processing. This is a challenging inference task with a complex symmetrical and multimodal posterior distribution. The model is provided in Supplementary Section 6. The observation is a vector of length 10 and $\bm{\theta}$ has dimension 2. SeMPLE is run with $K=40$, a value that is selected via BIC in 4.5 minutes. The total number of model simulations is set to $4\times 10^{4}$. C2ST and runtime results are reported in Figure 3. An exemplary posterior is shown in Figure 4. SeMPLE shows the best performance according to C2ST, and its runtime is about half that of SNL, and is about 10 times more efficient than SNPE-C. Adding the BIC overhead does not change the conclusion that SeMPLE is much more efficient. The runtime comparison with SNL is not particularly relevant, since the latter does not generally manage to explore the multimodal posterior (see Figure 4 and the posteriors in Supplementary Section 6).
@@
Figure 4: Hyperboloid. An example of marginal posteriors and samples from the last iteration of each algorithm. The SNL marginal posteriors are not reported as their inference is completely off.
@@ 5.3 Lotka-Volterra
The Lotka-Volterra predator prey model is a classical example of a population dynamics model. We consider the stochastic Markov jump process version of the model, which is defined in Supplementary Section 9. This model, although being relatively simple and characterized by a set of three interactions on the two species, each with a corresponding rate $\theta_{j}$ ($j=1,2,3$), encompasses many of the difficulties associated with larger, more complex systems. The observations consider the addition of non-negligible measurement error to the simulated dynamics, and the standard deviation $\sigma$ of this noise term is inferred as an additional parameter, resulting in a total of four model parameters to infer $(\theta_{1},\theta_{2},\theta_{3},\sigma)$. For this model, we are not primarily focusing on the performance comparison to other neural-based methods, as this version of the model, as a Markov jump process, is not included in SBIBM. Moreover, running times for a single model call are random even for repeated calls using the same parameter value. We use the Gillespie algorithm as detailed in Supplementary Section 9, and the number of “reactions” at any given $\bm{\theta}$ is stochastic. Due to the reasons above, the primary focus is to show that high quality inference can be obtained with SeMPLE using a comparably small computational effort, and compare our results against two versions of sequential Monte Carlo ABC (SMC-ABC). A “reference posterior” is obtained with an efficient version of SMC-ABC called blockedopt in Picchini and Tamborrino (2022), which is run for very many model simulations, around 3.3 million, see the Supplementary Material for details. We also compare it with a more standard off-the-shelf SMC-ABC named standard. SeMPLE inference is obtained with only 30,000 model simulations. In Figure 5 we also include the corresponding inference obtained with around 30,000 model simulations for both blockedopt and standard SMC-ABC. With only 30,000 model simulations, SeMPLE manages to produce inference for all parameters similar to the inference returned after 3.3 million model simulations from blockedopt SMC-ABC, a 100-fold computational saving. We observe that, unlike SMC-ABC, SeMPLE produces accurate posterior inference for all model parameters, including the noise parameter $\sigma$, but with a much smaller number of model simulations compared to SMC-ABC. The latter fact, namely the SMC-ABC inefficiency in terms of model simulations, was not unexpected and known from Greenberg et al. (2019); Lueckmann et al. (2021). However it is reassuring to notice the excellent inference quality brought by SeMPLE in this challenging case study.

### 16 Mar 2024 v2

@@ 4 SeMPLE: sequential learning of likelihoods and posteriors
We have shown how GLLiM can provide amortized estimations of both the likelihood and the posterior distribution at no additional cost, given training data. This GLLiM feature is exploited in our proposed Sequential Mixture Posterior and Likelihood Estimation (SeMPLE), to obtain a sequence of increasingly improved samplers, to more accurately and more efficiently sample from the targeted posterior of observed data $\bm{y}_{o}$. Similarly to SNL (Papamakarios et al., 2019), given a likelihood approximation $q_{\tilde{\bm{\phi}}}(\bm{y}_{o}\,|\,\bm{\theta})$, SeMPLE targets an approximate posterior $\tilde{p}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ with posterior sampling being carried out using a Markov Chain Monte Carlo (MCMC) procedure. Typically, the MCMC scheme used in SNL is slice sampling (Neal, 2003), which has the advantage to be tuning-free but is likely to have difficulties in exploring high dimensional posteriors with complex multimodalities. As a matter of fact, Papamakarios et al. (2019) recommend SNL “at most for tens of parameters that do not have pathologically strong correlations”. In SeMPLE, since GLLiM is able to learn a closed-form approximation of the posterior distribution (in addition to a likelihood function), such posterior is used as a proposal distribution within a Metropolis-Hastings (MH) algorithm. SeMPLE, by using an efficient EM procedure to fit a flexible proposal distribution, is expected to benefit from a MH step that better guides simulations towards more likely parameter values. To ease the comparison with SNL, the latter is recalled in Appendix D, while SeMPLE is summarized in Algorithms 1-2.
@@
Algorithm 1 shows that the targeted approximate posterior for $\bm{\theta}$ is a function denoted with $\hat{p}_{r}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ which is sequentially obtained, where $r=0,...,R$ denotes the SeMPLE rounds for the current approximation [^2] We use the term round instead of the most common iteration, as the latter could be confused with the MCMC iterations that are run within a given SeMPLE round.. In the following we denote by $\bm{\theta}^{(r)}$ and $\bm{y}^{(r)}$ a parameter and an observation simulated in round $r$, respectively. At first ($r=0$) $\hat{p}_{0}(\bm{\theta})$ is initialised at the prior $p(\bm{\theta})$. An initial batch of $N$ pairs $\mathcal{D}_{0}=\{\bm{\theta}_{n}^{(0)},\bm{y}_{n}^{(0)}\}_{n=1}^{N}$ produced from the prior-predictive distribution (step 3) constitutes the starting training data for GLLiM, and the parameters $\tilde{\phi}_{0}$ and ${\phi}_{0}$ are obtained via EM within GLLiM to produce the initial surrogate likelihood and posterior $q_{\tilde{\bm{\phi}}_{0}}(\bm{y}\,|\,\bm{\theta})$ and $q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y})$ respectively (steps 8-9). The initial surrogate posterior becomes the current target (step 10) $\tilde{p}_{1}(\bm{\theta})\equiv q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y}_{o})$. In the next round ($r=1$), the latter can be sampled from without MH (step 3) since $\tilde{p}_{1}(\bm{\theta})$ is defined as a Gaussian mixture, from which it is trivial to sample $N$ parameters $\{\bm{\theta}_{n}^{(1)}\}_{n=1}^{N}$. These $N$ draws are then plugged into the model simulator to produce $\bm{y}_{n}^{(1)}\sim p(\bm{y}\,|\,\bm{\theta}_{n}^{(1)})$ and define $\mathcal{D}_{1}=\{\bm{\theta}_{n}^{(1)},\bm{y}_{n}^{(1)}\}_{n=1}^{N}$ as the new training data (we discard $\mathcal{D}_{0}$ for future rounds as it is deemed too uninformative for training). GLLiM is then fitted to $\mathcal{D}_{1}$ and the corresponding surrogate likelihood and surrogate posterior are obtained. From this step on, the surrogate posterior is employed as a proposal distribution within a tuning-free MH algorithm (Algorithm 2) since, for an arbitrary prior $p(\bm{\theta})$, when $r\geq 2$ we do not explicitly know which distribution is proportional to the product $p(\bm{\theta})q_{\tilde{\phi}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})$ (for $r=1$ we know by construction that the targeted posterior is a Gaussian mixture), except for the special case of Gaussian and Uniform priors (see Papamakarios & Murray, 2016). In Algorithm 2, the acceptance probability of a proposed $\bm{\theta}^{*}$ is (for $r>1$)
@@
Importantly, our MCMC procedure is tuning-free. We use an independence sampler (Robert & Casella, 2004), relying on GLLiM to provide an informative proposal function. This proposal function is trained on draws accepted in previous rounds which are, reasonably, more spread than the currently targeted posterior. In addition, being a mixture model, it is suitable to deal with multimodalities in the target. This greatly simplifies the construction of SeMPLE as no complex tuning is necessary for the MCMC sampler. While $K$ could in principle be set arbitrarily large, in practice this may slow the EM algorithm convergence. In SeMPLE, the value of $K$ is allowed to decrease between rounds in the update $K$ step lines 8 and 13 in Algorithm 1. Mixture components having probabilities $\tilde{\eta}_{k}$ smaller than a threshold are discarded. In most cases we find that setting a starting value of $K=20$ or 30 produces good results, with a threshold for $\tilde{\eta}_{k}$ typically set to 0.005 (multiple hyperboloid, Ornstein-Uhlenbeck, Lotka-Volterra models), 0.03 (Bernoulli GLM model) or simply 0 for no deletion (Two Moons model). This is to prevent having too many irrelevant components, that could be problematic when fitting the mixture to training data produced via MCMC, as the latter may not manage to visit tiny modes. It is also possible to choose the starting value of $K$ in a principled way. In Deleforge et al. (2014), the Bayesian Information Criterion (BIC) is used to select $K$, as we do in our examples, see Appendix B. An appropriate value for $K$ can be learned by repeatedly computing the BIC on realizations of the prior-predictive distribution, before SeMPLE is started. This is an amortized procedure, as the selected values can be reused regardless the specific observed data $\bm{y}_{o}$. The impact on the overall computational budget is negligible. We have also experimented with a version of SeMPLE that learns a corrected posterior (see the optional part in line 16 of Algorithm 1) using the correction formula from Papamakarios & Murray (2016). However, we found that learning a surrogate likelihood was much more beneficial, providing better inference coupled with a higher acceptance rate. Notice that none of our reported results use the optional corrected-posterior approach.
@@ 5 Numerical illustrations
We illustrate SeMPLE on several simulation studies, see also the Appendices for extra results and additional models. SeMPLE is written in R and uses the GLLiM implementation from the xLLiM package (Perthame et al., 2022). Comparisons with neural-based likelihood and posterior estimation methods (SNL and SNPE-C) use the Python implementations in the SBIBM package for SBI benchmarking (Lueckmann et al., 2021), and architecture specifications for SBIBM are in Appendix A. To compare approximate posteriors with a suitable reference posterior, we consider the classifier two-samples test (C2ST), as implemented in SBIBM, and Wasserstein distances via the POT Python package (Flamary et al., 2021). C2ST varies in the interval [0.5,1], the lower the better. It equals 0.5 when samples from two distributions appear statistically indistinguishable and equals 1 when the classifier is able to perfectly separate samples as belonging to two different distributions. We also report runtimes in figures and resource requirements in Table 1. For SeMPLE, the time required to select $K$ via BIC is reported separately and not taken into account in the figures below. For some case studies, such as Lotka-Volterra, we merely picked a starting value of $K$ without running the BIC procedure. Adding this overhead time does not change the overall conclusion that SeMPLE is much more time-efficient than the other methods. Moreover, the selection of $K$ is an amortized procedure, hence it does not need to be repeated for different observations of the same model.
@@
In the next two sections 5.1 and 5.2, we illustrate results from two models having multimodal posteriors, in section 5.3 we summarise the Lotka-Volterra experiment, and in section 5.5 and Table 2 we summarise results from all other experiments detailed in the Appendices. For comparison, all methods are given the same budget of model simulations. For SNL and SNPE-C the SBIBM default of distributing these simulations uniformly across $R=10$ rounds is used, but to ease comparison we also run SNL and SNPE-C with the same number of rounds used for SeMPLE, which is $R=4$, except for the Bernoulli GLM model where $R=2$ suffices. In fact, we observe that SeMPLE does not need as many sequential steps as the other methods. It may be argued that SNL could perform better if a larger number of MCMC iterations was used, at the cost of a smaller number of rounds $R$. Results in the Appendices show that both SNL and SNPE-C perform worse under this alternative configuration. Regarding SeMPLE, when the computational budget limits the number of model simulations, we found it beneficial to keep the number of rounds $R$ between 2 and 3, where $R=2$ implies that we have exactly one round where MCMC is used, and this is typically already enough for satisfying inference, though usually an additional round with MCMC (hence $R=3$) produces further refined inference. In our work we show results up to $R=4$, but in general, four rounds are not recommended. The algorithm performance relies first on the GLLiM ability to learn reasonably accurate likelihood and posterior approximations and then on the MCMC sampler to explore the parameter space. Both steps require a large enough number of model simulations, which depends mainly on the model dimension.
@@ 5.1 Two Moons
The Two Moons model is a standard benchmark example in SBI (Greenberg et al., 2019; Lueckmann et al., 2021; Picchini & Tamborrino, 2022) to illustrate how algorithms deal with multimodality and the local crescent shapes of the posterior distribution. The model itself is defined in Appendix E. Our experiments are run on the same 10 datasets provided in SBIBM, each of them being a vector $\bm{y}_{0}$ of length two. We use $K=30$, a value we determine as described in Appendix B (this determination only required 90 seconds). We run SeMPLE, SNL and SNPE-C algorithms for a total budget of $10^{4}$ model simulations, distributed across inference rouces, and we tested both $R=4$ and $R=10$ rounds. The obtained samples are then compared with the corresponding posterior reference samples (obtained via MCMC, since the likelihood is tractable) provided in SBIBM. The performance metric results are averaged over the 10 different data sets. Figure 2 shows C2ST and runtimes values. SeMPLE provides the best approximation while SNL performs the worst. The comparison with SNL is particularly interesting as while both SNL and SeMPLE incorporate a tuning-free MCMC step, SeMPLE is able to visit both modes with an acceptance rate of 60-70%, while SNL often struggles in efficiently exploring a multimodal posterior as demonstrated by the higher C2ST values. Moreover, in this simple example the runtime for SeMPLE is about half the SNPE-C runtime, and is about three times lower than the SNL runtime. In other examples, (e.g. Figure 3), the speedup provided by SeMPLE is even more dramatic. All these benefits from SeMPLE come with a very small impact on the memory usage, see Section 5.4 for details.
@@ 5.2 Multiple hyperboloid model
The multiple hyperboloid model (Forbes et al., 2022) is constructed from a sound source localization problem in audio processing. This is a challenging inference task with a complex symmetrical and multimodal posterior distribution. The model is provided in Appendix F. The observation is a vector of length 10 and $\bm{\theta}$ has dimension 2. SeMPLE is run with $K=40$, a value that is selected via BIC in 4.5 minutes. The total number of model simulations is set to $4\times 10^{4}$. C2ST and runtime results are reported in Figure 3. An exemplary posterior is shown in Figure 4. SeMPLE shows the best performance according to C2ST, and its runtime is about half that of SNL, and is about 10 times more efficient than SNPE-C. Adding the BIC overhead does not change the conclusion that SeMPLE is much more efficient. The runtime comparison with SNL is not particularly relevant, since the latter does not generally manage to explore the multimodal posterior (see Figure 4 and the posteriors in Appendix F).
@@
Figure 4: Hyperboloid. An example of marginal posteriors and samples from the last round of each algorithm. The SNL marginal posteriors are not reported as their inference is completely off.
@@ 5.3 Lotka-Volterra
The Lotka-Volterra predator prey model is a classical example of a population dynamics model. We consider the stochastic Markov jump process version of the model, which is defined in Appendix I. This model, although being relatively simple and characterized by a set of three interactions on the two species, each with a corresponding rate $\theta_{j}$ ($j=1,2,3$), encompasses many of the difficulties associated with larger, more complex systems. The observations consider the addition of non-negligible measurement error to the simulated dynamics, and the standard deviation $\sigma$ of this noise term is inferred as an additional parameter, resulting in a total of four model parameters to infer $(\theta_{1},\theta_{2},\theta_{3},\sigma)$. For this model, we are not primarily focusing on the performance comparison to other neural-based methods, as this version of the model, as a Markov jump process, is not included in SBIBM. Moreover, running times for a single model call are random even for repeated calls using the same parameter value. We use the Gillespie algorithm, as explained in Appendix I, and the number of “reactions” at any given $\bm{\theta}$ is stochastic. Due to the reasons above, the primary focus is to show that high quality inference can be obtained with SeMPLE using a comparably small computational effort, and compare our results against two versions of sequential Monte Carlo ABC (SMC-ABC). A “reference posterior” is obtained with an efficient version of SMC-ABC called blockedopt in Picchini & Tamborrino (2022), which is run for very many model simulations, around 3.3 million, see Appendix I for details. We also compare it with a more standard off-the-shelf SMC-ABC named standard. SeMPLE inference is obtained with only 30,000 model simulations. In Figure 5 we also include the corresponding inference obtained with around 30,000 model simulations for both blockedopt and standard SMC-ABC. With only 30,000 model simulations, SeMPLE manages to produce inference for all parameters similar to the inference returned after 3.3 million model simulations from blockedopt SMC-ABC, a 100-fold computational saving. We observe that, unlike SMC-ABC, SeMPLE produces accurate posterior inference for all model parameters, including the noise parameter $\sigma$, but with a much smaller number of model simulations compared to SMC-ABC. The latter fact, namely the SMC-ABC inefficiency in terms of model simulations, was not unexpected and known from Greenberg et al. (2019); Lueckmann et al. (2021). However it is reassuring to notice the excellent inference quality brought by SeMPLE in this challenging case study.